OpenStudy (anonymous):

fin the lim as x approaches π/3 of 1/2-cosx over 2x-2π/3 help please!

OpenStudy (psymon):

\[\frac{ 1 }{ 2 }-\frac{ cosx }{ 2x }-\frac{ 2\pi }{ 3 }\]?

OpenStudy (anonymous):

fin the lim as x approaches π/3 of 1/2-cosx over 2x-2π/3

OpenStudy (psymon):

\[\frac{ \frac{ 1 }{ 2}-cosx }{ 2x-\frac{ 2\pi }{ 3 } }\] Just want to make sure I get it right.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my teacher told us the answer was square root of 3 over 4. but he never showed the work:/

OpenStudy (happinessbreaksbones):

psymon is so smart D:

OpenStudy (anonymous):

@happinessbreaksbones IKR!

OpenStudy (happinessbreaksbones):

he's amazing

OpenStudy (happinessbreaksbones):

and I'm glad he helped you :)

OpenStudy (anonymous):

yeah me too. or I woul've been staring at one problem for hours and still feeling blank.

OpenStudy (psymon):

Im trying to find the calc 1 way of doing this. I can do it the calc 2 way, haha

OpenStudy (anonymous):

the problem I posted?

OpenStudy (psymon):

Yeah. I can get the limit through a trick, but its not something you're supposed to know. Im multitasking, so I apologize for the wait.

OpenStudy (anonymous):

its ok.

OpenStudy (psymon):

Still trying to take a look. The answer you said before is right, just need to get it a different way, lol.

OpenStudy (anonymous):

lol my AP Cal teacher is crazy he gives us the answer and tells us to go home and work out the problem, and that we must know for the quizzes he's given out the next day

terenzreignz (terenzreignz):

Those fraction bars look really intimidating... can't there be just one? :3

OpenStudy (psymon):

That didnt see to do it, though, lol.

terenzreignz (terenzreignz):

\[\Large \frac{ \frac{ 1 }{ 2}-cosx }{ 2x-\frac{ 2\pi }{ 3 } }= \frac{\frac{1-2\cos(x)}{2}}{\frac{6x-2\pi}{3}}\]\[\Large = \frac{3-6\cos(x)}{12x -4\pi}\]

OpenStudy (psymon):

Still gives undefined there xD

OpenStudy (anonymous):

that's what I got!

OpenStudy (anonymous):

from there I stopped idnt know where to go

OpenStudy (psymon):

That x - pi is whats miserably getting in the way.

OpenStudy (anonymous):

for the numerator I did simplify and got 3(1-2cos)

OpenStudy (psymon):

Thats not the problem, its the x - pi. Thats what our focus needs to be on.

OpenStudy (anonymous):

yeah true

hartnn (hartnn):

i would say , just substitute y = x- pi/3 and simplify first...

OpenStudy (anonymous):

when you do that the answer is 0/0

hartnn (hartnn):

the answer will always be 0/0, unless we cancel out a common factor which makes num and denom=0

hartnn (hartnn):

[1/2 - cos (y+pi/3)]/(2y) easier to deal with ?

hartnn (hartnn):

expand cos (y+pi/3) now, can you ?

terenzreignz (terenzreignz):

What I did... combines elements of Master Hartnn's idea of substituting \[\Large y = x-\frac{\pi}3\]

terenzreignz (terenzreignz):

And the fact that \[\Large \lim_{p\rightarrow 0}\frac{1-\cos(p)}{p}=0\]

hartnn (hartnn):

and the fact that lim sin x/x = 0 when x->0

hartnn (hartnn):

i meant 1

terenzreignz (terenzreignz):

lol... me too slow ^_^

hartnn (hartnn):

@have_sabr , are you following us ? could you expand cos (y+pi/3) ? or did u get till that step?

OpenStudy (anonymous):

lol no I have to visualize it.

hartnn (hartnn):

which step are you now? (stuck at?)

terenzreignz (terenzreignz):

Why don't you start with this: \[\Large = \frac{3-6\cos(x)}{12x -4\pi}\] And set \[\large x = y+\frac{\pi}3\]

terenzreignz (terenzreignz):

And do the necessary substituting? By the way, \[\Large x \rightarrow \frac \pi 3\]\[\implies\]\[\color{blue}{\Large y \rightarrow0}\]

OpenStudy (anonymous):

i did what @terenzreignz said set x equal to pi/3

terenzreignz (terenzreignz):

I never said that ^_^ I said set \[\LARGE x = \color{red}{y+}\frac \pi 3\]

OpenStudy (anonymous):

lol omg im to slow to do this over the enternet. i would actually have to see exactly.

hartnn (hartnn):

\(\large \frac{ \frac{ 1 }{ 2}-cosx }{ 2x-\frac{ 2\pi }{ 3 } }=\frac{ \frac{ 1 }{ 2}-cosx }{ 2(x-\frac{ \pi }{ 3 } )}=\dfrac{\dfrac{1}{2}-\cos(y+\pi/3)}{2y}\) when x = y+pi/3

hartnn (hartnn):

got that? now can you expand cos(y+pi/3) ?

OpenStudy (anonymous):

the equation was a picture before and its just whole bunch of brackets and number which i didn't get. so you guys thanks for the trouble you went though but i just cant get. I'll see if my teacher ould really explain.

hartnn (hartnn):

just refresh your page...

OpenStudy (psymon):

@have_sabr Well, we have a bit of it laid out for you. You can expand like hartnn was saying by using the sum of cosines formula.

OpenStudy (psymon):

 I just drew what hartnn previously posted. Maybe you can see it when drawn?